OpenStudy (lgg23):

Find the volume of the solid obtained by rotating the region bounded by the given curve about the specified axis. x^2+(y-5)^2=16 about the y axis. Volume = __?__

OpenStudy (amistre64):

so we have a donut pretty much

OpenStudy (amistre64):

radius of 4; and we are 5 away from the center so yeah; a donut

OpenStudy (amistre64):

shell method might be appicable

OpenStudy (amistre64):



OpenStudy (amistre64):

\[\int 2pix\ f(x)\ dx\] \[f(x)=\sqrt{16-x^2}+5\] \[4pi\int_{2}^{10}x\sqrt{16-x^2}+5x\ dx\]

OpenStudy (amistre64):

one idea that might work, and be simpler is this: